# PaCo-FR: Patch-Pixel Aligned End-to-End Codebook Learning for Facial Representation Pre-training

### v1

@@ 1 Introduction
Facial representation learning is fundamental to human-centric AI [6, 6, 65, 49, 24], supporting applications like face recognition [64], identity verification [42], expression analysis [3], and digital avatar animation [62]. The quality of facial representations directly affects the robustness, fairness, and scalability of downstream systems. As facial imagery proliferates, there is growing demand for representations that are accurate, generalizable, data-efficient, and deployment-friendly.
@@
Meanwhile, large-scale visual pre-training has become a dominant approach in computer vision [1], especially in unsupervised or weakly supervised settings. Methods like MoCo [26, 11, 21], SimCLR [10], and CLIP [40] show that proxy tasks such as contrastive learning or masked modeling can produce transferable features. However, these general models often underperform on face-related tasks due to a lack of alignment with the unique regularities and semantics of human faces, highlighting the need for scalable yet facially-aware pre-training strategies.
To address this, recent work has explored face-specific pre-training. FaRL [65] uses the large-scale LAION-FACE dataset to train a vision-language model aligned with facial semantics, while MCF [56] applies contrastive learning to a 2M-image subset, achieving strong results. Despite these advances, further work is needed to fully exploit the spatial and semantic structures unique to faces.
Fully capturing these properties remains a key challenge in advancing facial representation learning. To address this, we conduct a detailed analysis of the characteristics present in facial data. As shown in Figure LABEL:fig:motivate(a), alignment procedures reveal substantial content similarities at corresponding facial positions, emphasizing the critical role of spatial coherence in facial representations. This motivates the use of masked image modeling (MIM) to enhance the model’s understanding of the inherent spatial structure of faces. Further analysis in Figure LABEL:fig:motivate(b) demonstrates how primary facial categories, such as eyes and mouths, become differentiated into nuanced subcategories influenced by attributes like makeup and expression. This insight underscores the importance of capturing fine-grained variations in facial features, which is essential for accurately distinguishing individuals and handling diverse facial variations.
@@
Building upon these observations, we propose PaCo-FR, an efficient unsupervised Facial Representation pre-training framework that advances the state-of-the-art by leveraging Patch-pixel alignment and end-to-end Codebook learning. Drawing inspiration from codebook-based generative models such as VQ-VAE and BEiT, our method integrates discrete token learning into facial modeling through a novel masking and replacement mechanism. Unlike conventional approaches, which treat image patches as independent entities, our method ensures spatial consistency by performing facial alignment, thereby maintaining the structural integrity of the face. After aligning the face, we divide the image into semantically coherent patches, each of which is assigned multiple candidate tokens within a codebook. These tokens are dynamically selected via a lightweight Belief Predictor, which replaces the original image patches with corresponding tokens. The model then reconstructs the masked image in an end-to-end manner, forcing it to learn both discriminative semantics and to preserve the geometric structure of facial features.
To enhance facial awareness further, we curate LAION-FACE-2M-crop, a high-quality subset of 2 million aligned facial images, obtained by applying landmark-guided cropping to the original LAION-FACE dataset. This carefully curated subset serves as the foundation for our pre-training pipeline, ensuring that the model is exposed to high-quality, spatially consistent facial data. We evaluate our pre-trained model across a variety of downstream tasks related to face recognition, attribute prediction, and other face-centric applications. Despite the relatively modest size of our training dataset, our method consistently outperforms existing approaches, demonstrating superior generalization and transferability. These results underscore the efficacy of our framework in capturing the fine-grained spatial and semantic features of faces, ultimately setting a new benchmark for facial representation learning.
In summary, our major contributions are as follows: (1) We propose a novel pre-training strategy that places the codebook at the image decoding end, resolving gradient back-propagation issues and unifying traditional two-stage codebook-based pre-training into a single-stage framework. (2) We introduce the Belief Predictor to inject attribute-aware priors into token selection, enhancing the semantic expressiveness and discriminability of the learned codebook. (3) Our end-to-end training enables direct patch-level token learning, improving the model’s capacity to capture facial structural and semantic patterns.
@@ 4.1 Experimental Settings
Method Pre-train $\mathrm{NME}_{\mathrm{inter\textit{-}ocular}}$ $\downarrow$ Full Pose Expr. Illum. MakeUp Occl. Blur Previous methods ESR [7] 11.13 25.88 11.47 10.49 11.05 13.75 12.20 SDM [61] 10.29 24.10 11.45 9.32 9.38 13.03 11.28 CFSS [66] 9.07 21.36 10.09 8.30 8.74 11.76 9.96 DVLN [58] 6.08 11.54 6.78 5.73 5.98 7.33 6.88 LAB [59] 5.27 10.24 5.51 5.23 5.15 6.79 6.12 Wing [23] 5.11 8.75 5.36 4.93 5.41 6.37 5.81 DeCaFA [13] 4.62 8.11 4.65 4.41 4.63 5.74 5.38 AWing [55] 4.36 7.38 4.58 4.32 4.27 5.19 4.96 ADNet [28] 4.14 6.96 4.38 4.09 4.05 5.06 4.79 Pre-trained methods Scratch None 4.95 8.95 5.22 4.84 5.19 6.31 5.46 FaRL [65] 20M 4.38 7.60 4.66 4.19 4.30 5.44 4.98 MCF [56] 2M 4.51 7.93 4.73 4.44 4.53 5.63 5.17 PaCo-FR 2M 4.35 7.50 4.52 4.22 4.18 5.43 5.02 FaRLft [65] 20M 4.03 6.81 4.32 3.92 3.87 4.70 4.54 MCFft [56] 2M 4.16 7.15 4.40 4.05 4.09 4.99 4.66 PaCo-FRft 2M 3.99 6.75 4.21 3.87 3.80 4.71 4.47

### v2

@@ 1 Introduction
Figure 1: (a) The pipeline of FaRL [52], (b) The pipeline of MCF [46], and (c) our proposed PaCo-FR. PaCo-FR is specifically designed for facial characteristics and optimize it as a learning objective for fatial representation pre-training.
@@
Facial representation learning is fundamental to human-centric AI [5, 52, 40, 19], supporting key applications such as face recognition [51], identity verification [34], expression analysis [3], and avatar animation [50]. The accuracy of facial representations crucially affects the robustness and fairness of downstream systems. With facial imagery proliferating in settings like social media, there is an urgent need for representations that are accurate, generalizable, and efficient for deployment.
@@
Meanwhile, large-scale visual pre-training has become the standard in computer vision [1], aided by methods like MoCo [22, 9, 17], SimCLR [8], and CLIP [32]. These approaches excel in extracting general visual features using unsupervised or weakly supervised learning on massive datasets. However, when applied to facial tasks, such models often fall short because they lack inductive alignment with the unique structures and semantics of human faces, pointing to the need for pre-training strategies tailored to facial data.
Recent research has started to address this gap with domain-specific approaches. FaRL [52] introduced a large face-centric dataset (LAION-FACE) and vision-language pre-training, while MCF [46] applied contrastive learning on a curated subset, showing the benefits of face-specific strategies. Nonetheless, fully exploiting the spatial regularities and fine-grained semantic details of faces remains a challenge.
To tackle this, we analyze key characteristics of facial data. As illustrated in Figure 2(a), faces exhibit strong spatial coherence, suggesting that masked image modeling can help capture their inherent structure. Figure 2(b) further shows that primary facial features, such as eyes and mouths, include fine-grained subcategories shaped by attributes like makeup and expression. Capturing these subtle variations is critical for distinguishing individuals and handling diverse facial appearances.
(a) Face alignment.
(b) Feature sub-categorization.
Figure 2: Two key phenomena observed in facial image analysis. (a) After face alignment, various facial elements can be matched to corresponding positions, enhancing consistency. (b) Facial features, such as eyes, are clustered into subcategories based on attributes like makeup and state, enabling finer classification.
@@
Based on the above insights, we present PaCo-FR: an efficient unsupervised facial representation pre-training framework that advances state-of-the-art performance through patch-pixel alignment and end-to-end codebook learning. Inspired by codebook-based models like VQ-VAE and BEiT, PaCo-FR incorporates discrete token learning into facial modeling using a novel masking and replacement strategy. Unlike conventional methods that treat image patches independently, PaCo-FR performs facial alignment to preserve spatial and structural integrity. The aligned image is divided into semantically meaningful patches, each mapped to candidate tokens from a codebook. A lightweight Belief Predictor dynamically selects tokens to replace the original patches, and the model reconstructs the masked image, learning both the semantic and geometric structure of facial features.
For high-quality training data, we curate LAION-FACE-2M-crop, a set of 2 million aligned facial images from LAION-FACE via landmark-based cropping. This ensures the model is pre-trained with spatially consistent facial data. Our approach is evaluated on face recognition, attribute prediction, and related tasks, consistently surpassing existing methods—demonstrating strong generalization and transfer abilities, even with a modest training set. These results highlight PaCo-FR’s effectiveness in capturing fine-grained spatial and semantic facial characteristics, establishing a new benchmark in facial representation learning. In summary, our main contributions are:
  • A new pre-training strategy that places the codebook at the decoding end, enabling end-to-end training and resolving back-propagation challenges of traditional two-stage frameworks;
  • The introduction of a Belief Predictor to inject attribute-aware priors into token selection, improving codebook expressiveness and discrimination;
  • End-to-end patch-level token learning for improved modeling of facial structural and semantic patterns.
@@ 4.1 Experimental Settings
Method Pre-train $\mathrm{NME}_{\mathrm{inter\textit{-}ocular}}$ $\downarrow$ Full Pose Expr. Illum. MakeUp Occl. Blur Previous methods LAB [48] 5.27 10.24 5.51 5.23 5.15 6.79 6.12 Wing [18] 5.11 8.75 5.36 4.93 5.41 6.37 5.81 DeCaFA [11] 4.62 8.11 4.65 4.41 4.63 5.74 5.38 AWing [45] 4.36 7.38 4.58 4.32 4.27 5.19 4.96 ADNet [23] 4.14 6.96 4.38 4.09 4.05 5.06 4.79 Pre-trained methods Scratch None 4.95 8.95 5.22 4.84 5.19 6.31 5.46 FaRL [52] 20M 4.38 7.60 4.66 4.19 4.30 5.44 4.98 MCF [46] 2M 4.51 7.93 4.73 4.44 4.53 5.63 5.17 PaCo-FR 2M 4.35 7.50 4.52 4.22 4.18 5.43 5.02 FaRLft [52] 20M 4.03 6.81 4.32 3.92 3.87 4.70 4.54 MCFft [46] 2M 4.16 7.15 4.40 4.05 4.09 4.99 4.66 PaCo-FRft 2M 3.99 6.75 4.21 3.87 3.80 4.71 4.47